Light of wavelength 700 nm is shown upon a double slit. The bright fringes are observed to be 2.1 mm apart on the viewing screen. Find the fringe spacing if the light were changed to a wavelength of 450 nm.
Answer #1

Distance between consecutive bright fringes S= wavelength*D/d

where D = distance of the screen

d = distance between slits

=> S proportional to wavelength

=> S2/S1 = wavelength2 /wavelength1

=> fringe spacing for wavelength = 450 nm is S2 = 450*2.1/700 = 1.35 mm